# DragPoser: Motion Reconstruction from Variable Sparse Tracking Signals via Latent Space Optimization

### v1

@@ 4.2. Overview
4.3. Architecture
Figure 2. Pose Autoencoder architecture for creating a structured latent space representing valid poses. Following the Variational Autoencoder approach, we train the network, beginning with the input pose data represented in dual quaternions to capture joint rotations and translations in root space. The Probabilistic Encoder $\mathcal{E}$ estimates the mean and variance parameters of the Gaussian distribution for the input pose. Subsequently, a latent vector $\mathbf{z}$ is sampled and passed into the Probabilistic Decoder $\mathcal{D}$ for pose reconstruction. Training involves two reconstruction losses: $\mathcal{L}_{Q}$ and $\mathcal{L}_{FK}$, in addition to $\mathcal{L}_{KLD}$ and $\mathcal{L}_{C}$ which structure the latent space, ensuring standard Gaussian alignment and latent space continuity, respectively.
@@ 4.3. Architecture
Our motion synthesis method relies on an optimization process in a latent space (explained in Section 4.4). The Pose Autoencoder component is used to construct this latent space from motion capture data. Autoencoders rely on the idea of having a bottleneck between two networks—Encoder and Decoder—which allows the networks to effectively project the supplied data to a space with fewer dimensions, known as latent space. However, this latent space is not explicitly structured and, thus, may have regions with invalid points. To address this, we based our Pose Autoencoder on the Variational Autoencoder (VAE) architecture (see Section 3), which forces the latent space to represent valid points in a specific domain. To design the Encoder and Decoder networks, we require a network capable of accurately reconstructing a large amount of motion capture data. We build on top of the work by Ponton et al. (2023) which proposes an autoencoder structure, with no explicit structure on the latent space, that uses skeleton-aware operations as building blocks (Aberman et al., 2020). Next, we introduce all modifications needed to construct our Pose Autoencoder; please refer to Ponton et al. (2023) for details on the skeleton-aware operations.
Probabilistic Encoder $\mathcal{E}$
Our Pose Autoencoder reconstructs single poses $\mathbf{x}^{(i)}\in\mathbb{R}^{J\times 4+3}$ as detailed in Section 4.1. However, the Encoder $\mathcal{E}$ input is amplified with each joint’s root space translations. This information helps the Encoder understand human motion better and explicitly encodes information about the dimensions of the skeleton. For this, we use dual quaternions (Andreou et al., 2022), which provide a unified and compact representation encoding both the rotational and translation information in orthogonal quaternions. Note that we encode the root’s world displacement in its corresponding dual quaternion. Thus, we define the function $DQ:\mathbb{R}^{J\times 4+3}\mapsto\mathbb{R}^{J\times 8}$, converting our original representation to dual quaternions.
@@
(4) $\mathbf{z}^{(i)}=\mu_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)+\mathbf{I}\,\sigma_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)\epsilon$
@@
(5) $\mathbf{\hat{x}}^{(i)}=\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{z}^{(i)})$
Reconstruction Loss $\mathcal{L}_{Q}$
@@
(6) $\mathcal{L}_{Q}=MSE(\mathbf{x}^{(i)},\mathbf{\hat{x}}^{(i)})$
Forward Kinematics Loss $\mathcal{L}_{FK}$
To facilitate the neural network’s understanding of the joint hierarchy, we additionally include an FK-based loss (Pavllo et al., 2018, 2020). We compute the global position of each joint using FK and compare it with the ground truth. Thus, we define the function $FK:\mathbb{R}^{J\times 4+3}\mapsto\mathbb{R}^{J\times 3}$, and the loss as follows:
(7) $\mathcal{L}_{FK}=MSE\left(FK(\mathbf{x}^{(i)}),FK(\mathbf{\hat{x}}^{(i)})\right)$
KL Divergence Loss $\mathcal{L}_{KLD}$
@@
(8) $\mathcal{L}_{KLD}$ $=\langle{1+\log\left(\sigma^{(i)}\right)-\mu^{(i)}-\sigma^{(i)},\mathbf{1}}\rangle$
(9) $\mu^{(i)}$ $=\mu_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)$
(10) $\sigma^{(i)}$ $=\sigma_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)$
Continuity Loss $\mathcal{L}_{C}$
@@
(11) $\mathcal{L}_{C}=MSE\left(\mathbf{x}^{(i+1)},\mathcal{D}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}^{(i+1)})\right)$
(12) $\hat{\mathbf{z}}^{(i+1)}=\mathbf{z}^{(i)}-\nabla_{\mathbf{z}}MSE(\hat{\mathbf{x}}^{(i)},\mathbf{x}^{(i+1)})$
(13) $\nabla_{\mathbf{z}}MSE(\hat{\mathbf{x}}^{(i)},\mathbf{x}^{(i+1)})=\mathbf{J}^{\mathrm{T}}_{\mathbf{z}}\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{z}^{(i)})\cdot 2(\mathbf{x}^{(i+1)}-\hat{\mathbf{x}}^{(i)})$
Figure 3 illustrates the role of the Continuity Loss ($\mathcal{L}_{C}$) in structuring the latent space based on the temporal sequence of poses. In the absence of $\mathcal{L}_{C}$, the Pose Autoencoder clusters similar poses without considering their temporal sequence, which may result in an undesirable pose being traversed between two consecutive poses during optimization. This can cause the optimization process to become trapped in local minima. Conversely, incorporating $\mathcal{L}_{C}$ encourages the latent space to organize itself by pose similarity and temporal succession, ensuring that one pose leads smoothly to the next. This structure significantly reduces the risk of encountering local minima and facilitates a more reliable pose transition within the optimization step.
Figure 3. Demonstration of how the continuity loss $\mathcal{L}_{C}$ structures the latent space to reflect the temporality of data. The left diagram shows a latent space learned without $\mathcal{L}_{C}$, where similar poses are clustered together, but transitioning smoothly from one pose to the next is not guaranteed—potentially resulting in getting stuck in local minima. The right diagram illustrates the latent space with $\mathcal{L}_{C}$, encouraging consecutive poses from the original motion sequence to be within reach of a single optimization step. Notice the ability to transition from the last green pose to either the first red or blue pose, representing two potential future sequences. Each pose is represented as a 2D Gaussian distribution, while the mean predicted by the Encoder is represented with a symbol.
@@
The Pose Autoencoder effectively constructs a structured latent space for decoding single poses. However, this space lacks two critical properties. First, it is unbounded, assuming arbitrary shapes as combinations of Gaussian distributions. Second, while the Continuity Loss promotes poses to be reachable during the optimization step, it remains unclear which pose to select for temporal coherence among all possible reachable poses. This implies a lack of a clear understanding of the sequential progression of poses during optimization. To address these issues, we propose training a neural network—Temporal Predictor—to predict the most likely sequences of latent vectors based on their coherence with previous sequences. These predictions can then inform and guide the Pose Optimizer process. This approach not only helps navigate away from invalid regions in the latent space but also provides the system with an understanding of the temporal sequence of the data. A visual overview of the Temporal Predictor is depicted in Figure 4.
To maximize the span of previous and future latent vector sequences, without significantly enlarging the network, the Temporal Predictor $\mathcal{T}$ operates at a reduced frequency. Specifically, $\mathcal{T}$ is executed every $n$ frames, with $n=4$ in our experiments. We previously used $i$ to denote frames; now, we introduce a new index $j=\lceil i/n\rceil$. For indexing standard latent vectors $\mathbf{z}^{(i)}$, our frame rate is set to 60 fps, which corresponds to the standard frame rate of the Pose Optimizer. For indexing predicted latent vectors by the Temporal Predictor $\mathbf{z_{t}}^{(j)}$, we use the lower frame rate. The Temporal Predictor $\mathcal{T}_{\boldsymbol{\psi}}$, with parameters $\boldsymbol{\psi}$, adopts the Transformer architecture, comprising Encoder ($\mathcal{T}^{E}_{\boldsymbol{\psi}}$) and Decoder ($\mathcal{T}^{D}_{\boldsymbol{\psi}}$) models.
@@
The Encoder’s role is to contextualize the Temporal Predictor’s latent vector prediction. The iterative prediction of latent vectors in the Decoder, however, addresses different scenarios. For sparse inputs with minimal ambiguity, like six sensors (one per limb and the root), we set $W_{f}=1$. Therefore, the Temporal Predictor is only helping the Pose Optimizer process to stay within the distribution of valid poses. In cases of higher ambiguity, like four sensors (head, hands, and root), we increase $W_{f}$ to 16 or 60, letting the Decoder predict the next possible latent vectors. The Encoder handles actual past data and the Decoder makes future predictions. If the Decoder mainly gets data from the Encoder, we leave little space for prediction. Therefore, increasing $W_{f}$ effectively helps during the prediction process.
4.4. Pose Optimizer
Figure 4. Motion reconstruction process via latent space optimization. The procedure begins with an initial latent vector from which the pose is decoded. Constraints are then dynamically established as loss functions, and the gradients of these losses are computed via backpropagation within the latent space. The decoded pose is iteratively refined through this process, either until a maximum iteration count is reached or certain predefined metrics are satisfied. This optimization is guided by a Temporal Predictor network that employs a Transformer architecture. The Encoder module generates features from past poses, and the Decoder module predicts future poses. While the Pose Optimizer operates at every application frame, the Temporal Predictor is executed at every $n$ frames (with $n=4$ in our experiments).
The preceding sections detailed the components for creating a structured latent space to generate continuous and valid poses. However, using the Pose Autoencoder, we cannot directly synthesize full-body motion from sparse input. Taking inspiration from Pan et al. (2023), who demonstrate image modification in a GAN’s latent space through optimization, we crafted a similar latent space optimization process for pose prediction. This process is visually summarized in Figure 4 and outlined in Algorithm 1.
@@ 4.4. Pose Optimizer
Given a frame $i$, with its sparse input $\mathbf{s}^{(i)}$, and an initial pose in latent form $\mathbf{z}^{(i-1)}$, our goal is to predict the pose $\hat{\mathbf{x}}^{(i)}$ that best satisfies the constraints defined. We begin by decoding the initial latent vector to a pose using the Decoder of the Pose Autoencoder with fixed parameters: $\hat{\mathbf{x}}^{(i-1)}=\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{z}^{(i-1)})$. We then employ the sparse Forward Kinematics function $FK^{\prime}:\mathbb{R}^{J\times 4+3}\mapsto\mathbb{R}^{S\times(4+3)}$ to match the full-body pose with the global joint positions of the sparse input. The pose optimization loss, $\mathcal{L}_{PO}=MSE\left(FK^{\prime}(\hat{\mathbf{x}}^{(i)}),\mathbf{s}^{(i)}\right)$, encapsulates our constraints, which can be extended as needed, for instance, by adding or removing sensors on the fly, or imposing arbitrary limits on specific joints.
We backpropagate this loss through the Decoder to the latent space, obtaining a gradient to guide the update of $\mathbf{z}^{(i-1)}$ towards fulfilling the constraints. This process is similar to that of training the Decoder, but this time the parameters are fixed. The updated latent vector, $\hat{\mathbf{z}}^{(i)}=\mathbf{z}^{(i-1)}-\lambda_{PO}\nabla_{z}\mathcal{L}_{PO}$, is then decoded back to a pose $\hat{\mathbf{x}}^{(i)}=\mathcal{D}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}^{(i)})$. This iterative process is repeated until the positional and rotational error thresholds of the sparse tracking signals are satisfied or a maximum iteration count is reached.
Additionally, we integrate the Temporal Predictor’s output (see Section 4.3), $\mathbf{z_{t}}^{(j)}$, to avoid invalid latent space regions and to maintain temporal coherence, particularly when sensor input is sparse. This dual optimization, combining pose and temporal constraints, ensures our predictions remain both valid within the latent space and temporally consistent with the input data.
5. Experiments and Evaluation

### v2

@@ 4.2 Overview
4.3 Architecture
Figure 1: Pose Autoencoder architecture for creating a structured latent space representing valid poses. Following the Variational Autoencoder approach, we train the network, beginning with the input pose data represented in dual quaternions to capture joint rotations and translations in root space. The Probabilistic Encoder $\mathcal{E}$ estimates the mean and variance parameters of the Gaussian distribution for the input pose. Subsequently, a latent vector $\mathbf{z}$ is sampled and passed into the Probabilistic Decoder $\mathcal{D}$ for pose reconstruction. Training involves two reconstruction losses: $\mathcal{L}_{Q}$ and $\mathcal{L}_{FK}$, in addition to $\mathcal{L}_{KLD}$ and $\mathcal{L}_{C}$ which structure the latent space, ensuring standard Gaussian alignment and latent space continuity, respectively.
@@ 4.3 Architecture
Our motion synthesis method relies on an optimization process in a latent space (explained in Section 4.4). The Pose Autoencoder component is used to construct this latent space from motion capture data. Autoencoders rely on the idea of having a bottleneck between two networks—Encoder and Decoder—which allows the networks to effectively project the supplied data to a space with fewer dimensions, known as latent space. However, this latent space is not explicitly structured and, thus, may have regions with invalid points. To address this, we based our Pose Autoencoder on the Variational Autoencoder (VAE) architecture (see Section 3), which forces the latent space to represent valid points in a specific domain. To design the Encoder and Decoder networks, we require a network capable of accurately reconstructing a large amount of motion capture data. We build on top of the work by Ponton et al. [ponton2023sparseposer] which proposes an autoencoder structure, with no explicit structure on the latent space, that uses skeleton-aware operations as building blocks [Aberman:2020]. Next, we introduce all modifications needed to construct our Pose Autoencoder; please refer to Ponton et al. [ponton2023sparseposer] for details on the skeleton-aware operations.
Probabilistic Encoder
Our Pose Autoencoder reconstructs single poses $\mathbf{x}^{(i)}\in\mathbb{R}^{J\times 4+3}$ as detailed in Section 4.1. However, the input of the Encoder $\mathcal{E}$ is extended with each joint’s root space translations. This information helps the Encoder understand human motion better by explicitly encoding information about the dimensions of the skeleton. For this, we use dual quaternions [Andreou:2022], which provide a unified and compact representation encoding both the rotational and translation information in orthogonal quaternions, as used in previous work [ponton2023sparseposer]. We encode the root’s world displacement in its corresponding dual quaternion. Thus, we define the function $DQ:\mathbb{R}^{J\times 4+3}\mapsto\mathbb{R}^{J\times 8}$, converting our original representation to dual quaternions.
@@
$\mathbf{z}^{(i)}=\mu_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)+\mathbf{I}\,\sigma_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)\epsilon$ (4)
@@
$\mathbf{\hat{x}}^{(i)}=\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{z}^{(i)})$ (5)
Reconstruction Loss
@@
$\mathcal{L}_{Q}=MSE(\mathbf{x}^{(i)},\mathbf{\hat{x}}^{(i)})$ (6)
Forward Kinematics Loss
To facilitate the neural network’s understanding of the joint hierarchy, we additionally include an FK-based loss [Pavllo:2018, Pavllo:2020]. We compute the global position of each joint using FK and compare it with the ground truth. Thus, we define the function $FK:\mathbb{R}^{J\times 4+3}\mapsto\mathbb{R}^{J\times 3}$, and the loss as follows:
$\mathcal{L}_{FK}=MSE\left(FK(\mathbf{x}^{(i)}),FK(\mathbf{\hat{x}}^{(i)})\right)$ (7)
KL Divergence Loss
@@
$\mathcal{L}_{KLD}$ $=\frac{1}{2}\langle{1+\log\left((\sigma^{(i)})^{2}\right)-(\mu^{(i)})^{2}-(\sigma^{(i)})^{2},\mathbf{1}}\rangle$ (8)
$\mu^{(i)}$ $=\mu_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)$ (9)
$\sigma^{(i)}$ $=\sigma_{\boldsymbol{\phi}}\!\left(DQ(\mathbf{x}^{(i)})\right)$ (10)
Continuity Loss
@@
$\mathcal{L}_{C}=MSE\left(\mathbf{x}^{(i+1)},\mathcal{D}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}^{(i+1)})\right)$ (11)
$\hat{\mathbf{z}}^{(i+1)}=\mathbf{z}^{(i)}-\nabla_{\mathbf{z}}MSE(\hat{\mathbf{x}}^{(i)},\mathbf{x}^{(i+1)})$ (12)
$\nabla_{\mathbf{z}}MSE(\hat{\mathbf{x}}^{(i)},\mathbf{x}^{(i+1)})=\mathbf{J}^{\mathrm{T}}_{\mathbf{z}}\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{z}^{(i)})\cdot 2(\mathbf{x}^{(i+1)}-\hat{\mathbf{x}}^{(i)})$ (13)
where $\mathbf{J}^{\mathrm{T}}_{\mathbf{z}}$ is the transpose of the Jacobian matrix of the Decoder $\mathcal{D}$ with respect to $\mathbf{z}$. Figure 2 illustrates the role of the Continuity Loss ($\mathcal{L}_{C}$) in structuring the latent space based on the temporal sequence of poses. In the absence of $\mathcal{L}_{C}$, the Pose Autoencoder clusters similar poses without considering their temporal sequence, which may result in an undesirable pose being traversed between two consecutive poses during optimization. This can cause the optimization process to become trapped in local minima. Conversely, incorporating $\mathcal{L}_{C}$ encourages the latent space to organize by pose similarity and temporal succession, ensuring that one pose leads smoothly to the next. By mitigating the risk of local minima, this structure allows for more reliable pose transitions and faster optimization (demonstrated by the reduced iteration count in Section 5.4).
Figure 2: Demonstration of how the continuity loss $\mathcal{L}_{C}$ structures the latent space to reflect the temporality of data. The left diagram shows a latent space learned without $\mathcal{L}_{C}$, where similar poses are clustered together, but transitioning smoothly from one pose to the next is not guaranteed—potentially resulting in getting stuck in local minima. The right diagram illustrates the latent space with $\mathcal{L}_{C}$, encouraging consecutive poses from the original motion sequence to be within reach of a single optimization step. Notice the ability to transition from the last green pose to either the first red or blue pose, representing two potential future sequences. Each pose is represented as a 2D Gaussian distribution, while the mean predicted by the Encoder is represented with a symbol.
@@
The Pose Autoencoder effectively constructs a structured latent space for decoding single poses. However, this space lacks two critical properties. First, it is unbounded, assuming arbitrary shapes as combinations of Gaussian distributions. Second, while the Continuity Loss promotes poses to be reachable during the optimization step, it remains unclear which pose to select for temporal coherence among all possible reachable poses. This implies a lack of a clear understanding of the sequential progression of poses during optimization. We explored the idea of directly encoding sequences of poses in the latent vector. However, this resulted in a very slow optimization process and mostly non-accurate results due to the sparsity of the data. To address these issues, we propose training a neural network—Temporal Predictor—to predict the most likely sequences of latent vectors based on their coherence with previous sequences. These predictions can then inform and guide the Pose Optimizer process. This approach not only helps navigate away from invalid regions in the latent space but also provides the system with an understanding of the temporal sequence of the data. A visual overview of the Temporal Predictor is depicted in Figure 3.
To maximize the number of previous and future latent vector sequences, the Temporal Predictor $\mathcal{T}$ operates at a reduced frequency. Specifically, $\mathcal{T}$ is executed every $n$ frames, with $n=4$ in our experiments. We previously used $i$ to denote frames; now, we introduce a new index $j=\lceil i/n\rceil$. For indexing standard latent vectors $\mathbf{z}^{(i)}$, our frame rate is set to 60 fps, which corresponds to the standard frame rate of the Pose Optimizer. For indexing predicted latent vectors by the Temporal Predictor $\mathbf{z_{t}}^{(j)}$, we use the lower frame rate. The Temporal Predictor $\mathcal{T}_{\boldsymbol{\psi}}$, with parameters $\boldsymbol{\psi}$, adopts the Transformer architecture, comprising Encoder ($\mathcal{T}^{E}_{\boldsymbol{\psi}}$) and Decoder ($\mathcal{T}^{D}_{\boldsymbol{\psi}}$) models.
@@
The Encoder’s role is contextualizing the Temporal Predictor’s latent vector prediction. The iterative prediction of latent vectors in the Decoder, however, addresses different scenarios. For sparse inputs with minimal ambiguity, like six sensors (one per limb and the root), we set $W_{f}=1$. Therefore, the Temporal Predictor only helps the Pose Optimizer process to stay within the distribution of valid poses. In cases of higher ambiguity, like four sensors (head, hands, and root), we increase $W_{f}$ to 16 or 60, letting the Decoder predict the next possible latent vectors. The Encoder handles actual past data, and the Decoder makes future predictions. Therefore, increasing $W_{f}$ effectively helps during the prediction process.
The Temporal Predictor $\mathcal{T}$ is trained before the optimization process to learn about the data’s temporality. It is trained with a standard MSE loss function. Giving a motion clip from our database, we feed a sequence of poses to the Encoder $\mathcal{T}^{E}_{\boldsymbol{\psi}}$ and let the Decoder $\mathcal{T}^{D}_{\boldsymbol{\psi}}$ predict the following poses.
4.4 Pose Optimizer
Figure 3: Motion reconstruction process via latent space optimization. The procedure begins with an initial latent vector from which the pose is decoded. Constraints are then dynamically established as loss functions, and the gradients of these losses are computed via backpropagation within the latent space. The decoded pose is iteratively refined through this process, either until a maximum iteration count is reached or certain predefined metrics are satisfied. This optimization is guided by a Temporal Predictor network that employs a Transformer architecture. The Encoder module generates features from past poses, and the Decoder module predicts future poses. While the Pose Optimizer operates at every application frame, the Temporal Predictor is executed at every $n$ frames (with $n=4$ in our experiments).
The preceding sections detailed the components for creating a structured latent space to generate continuous and valid poses. However, using the Pose Autoencoder, we cannot directly synthesize full-body motion from sparse input. Taking inspiration from Pan et al. [pan:2023], who demonstrate image modification in a GAN’s latent space through optimization, we crafted a similar latent space optimization process for pose prediction. This process is visually summarized in Figure 3 and outlined in Algorithm 1.
@@ 4.4 Pose Optimizer
Given a frame $i$, with its sparse input $\mathbf{s}^{(i)}$, and an initial pose in latent form $\mathbf{z}^{(i-1)}$, our goal is to predict the pose $\hat{\mathbf{x}}^{(i)}$ that best satisfies the constraints defined. We begin by decoding the initial latent vector to a pose using the Decoder of the Pose Autoencoder with fixed parameters: $\hat{\mathbf{x}}^{(i-1)}=\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{z}^{(i-1)})$. We then employ the sparse Forward Kinematics function $FK^{\prime}:\mathbb{R}^{J\times 4+3}\mapsto\mathbb{R}^{S\times(4+3)}$ to match the full-body pose with the global joint positions and rotations of the sparse input. The pose optimization loss, $\mathcal{L}_{PO}=MSE\left(FK^{\prime}(\hat{\mathbf{x}}^{(i)}),\mathbf{s}^{(i)}\right)$, encapsulates our constraints, which can be extended as needed, for instance, by adding or removing sensors on the fly, or imposing arbitrary limits on specific joints, or constraining only positions or rotations.
We backpropagate this loss through the Decoder to the latent space, obtaining a gradient to guide the update of $\mathbf{z}^{(i-1)}$ towards fulfilling the constraints. This process is similar to that of training the Decoder, but this time the parameters are fixed. The updated latent vector, $\hat{\mathbf{z}}^{(i)}=\mathbf{z}^{(i-1)}-\lambda_{PO}\nabla_{z}\mathcal{L}_{PO}$, is then decoded back to a pose $\hat{\mathbf{x}}^{(i)}=\mathcal{D}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}^{(i)})$. This iterative process is repeated until the positional and rotational error thresholds of the sparse tracking signals are satisfied, a maximum iteration count is reached, or the gradient of the loss is very close to zero (e.g., $1\times 10^{-5}$). We found the latter condition to significantly speed up computations, as it is unlikely that any improvements will be made in this case.
Additionally, we integrate the Temporal Predictor’s output (see Section 4.3), $\mathbf{z_{t}}^{(j)}$, to avoid invalid latent space regions and to maintain temporal coherence, particularly when sensor input is sparse. This dual optimization, combining pose and temporal constraints, ensures our predictions remain both valid within the latent space and temporally consistent with the input data. The final latent vector update, as shown in Algorithm 1, incorporates both constraints. We maintain $\lambda_{PO}=1$ and adaptively adjust $\lambda_{T}$ based on the number of sensors used (0.15, 0.125, 0.02 for 3, 4, and 6 sensors, respectively).
5 Experiments and Evaluation